A man is equally likely to choose anyone of the three routes A, B, C from his house to a railway station, and his choice of route is not influenced by weather. If the weather is dry the probabilities of missing a train by routes A, B, C are 1/20, 1/10, 1/5 respectively. He sets out on a dry day and misses the train. a. Find the probability of missing the train . b. Find the probability of choosing route A and missing the train
1
Expert's answer
2021-05-26T03:02:34-0400

Let "P(A)" be the probability of using route A, "P(B)" be the probability of using route B, "P(C)" be the probability of using route C, "P(M)" be the probability of of missing a train.

Given "P(A)=P(B)=P(C)=\dfrac{1}{3}"

"P(M|A)=\dfrac{1}{20}, P(M|B)=\dfrac{1}{10}, P(M|C)=\dfrac{1}{5}"


a.

The Law of Total Probability


"P(M)=P(A)P(M|A)+P(B)P(M|B)"

"+P(C)P(M|C)=\dfrac{1}{3}(\dfrac{1}{20})+\dfrac{1}{3}(\dfrac{1}{10})+\dfrac{1}{5}(\dfrac{1}{3})"

"=\dfrac{1+2+4}{60}=\dfrac{7}{60}"

b. The Multiplication Rule


"P(C\cap M)=P(C)P(M|C)="

"=\dfrac{1}{3}(\dfrac{1}{20})=\dfrac{1}{60}"